OpenStudy (anonymous):

use L'Hôpital's rule to solve lim(x->0+) x^2x

OpenStudy (anonymous):

\[\lim_{x \rightarrow 0⁺}x²^x\] that's x to the power 2x

OpenStudy (anonymous):

so it's quite clearly of the form 0^0 and therefore it's ok to use L'Hôpital's rule

OpenStudy (anonymous):

but the 1st iteration gives 0^-1

OpenStudy (unklerhaukus):

\[\large\lim_{x \rightarrow 0^+}x^{2x} \]

OpenStudy (unklerhaukus):

\[\large\lim_{x \rightarrow 0^+}x^{2x}=\large\lim_{x \rightarrow 0^+}e^{\ln x^{2x}}=\large\lim_{x \rightarrow 0^+}e^{2x\ln x}=\]

Parth (parthkohli):

@UnkleRhaukus Darn.

OpenStudy (anonymous):

so then the derivative is lim(x->0) ln(x)*e^2x*ln(x)=ln(x)*x^2x

OpenStudy (unklerhaukus):

\[\large\lim_{x \rightarrow 0^+}e^{2x\ln x}=e^{~\lim\limits_{x \rightarrow 0^+}2x\ln x}=\]

Parth (parthkohli):

:')

OpenStudy (unklerhaukus):

\[\large e^{~\lim\limits_{x \rightarrow 0^+}2x\ln x}=e^{~2\lim\limits_{x \rightarrow 0^+}\frac{\ln x}{1/x}}=e^{~2\lim\limits_{x \rightarrow 0^+}\frac{1/x}{1/-x^2}}=e^{~2\lim\limits_{x \rightarrow 0^+}-x}\]

OpenStudy (anonymous):

so there is no need to take a derivative and the limit as x approaches 0 is 1?

OpenStudy (unklerhaukus):

i've used L'Hôp twice

OpenStudy (anonymous):

ok I see where you used L'Hôpital's. I'm still a little confused concerning some of the rules (1. why is it ok to put the limit to the exponent? 2. why is it ok to differentiate only the exponent? 3. What is the rule used in the 1st differentiation?). Thanks for the help anyways @UnkleRhaukus I'll try to wrap my head around it

OpenStudy (unklerhaukus):

1. the exponential function is continuos

OpenStudy (unklerhaukus):

2. \[{\lim\limits_{x \rightarrow 0^+}x\ln x}\]\[\qquad\qquad\qquad \text{let } {x=1/X}\]\[-{\lim\limits_{X \rightarrow \infty}\frac{\ln\tfrac 1X}{X}}\] \[-{\lim\limits_{X \rightarrow \infty}\frac{\tfrac 1X}{1}}\\-\lim\limits_{X \rightarrow \infty}\frac1X\\-\lim\limits_{x \rightarrow 0^+}x\]

OpenStudy (unklerhaukus):

make sense now?

OpenStudy (anonymous):

:D not exactly. I got point 1.

OpenStudy (anonymous):

3. Why does\[2x \ln(x)\] become\[2*\frac{ \ln(x) }{ 1/x }\] and not\[2*\ln(x)+\frac{ 1 }{ x }*2x=2*\ln(x)+2\] ?

OpenStudy (unklerhaukus):

well this is L'Hôp, we had the form \[m\times n \to (0)\times (-\infty)\]so we differentiate \(m\) and \(n\) separately (we are not using the product rule here)